# Observational constraints on the spin/anisotropy of the CCOs of Cassiopeia A, Vela Jr. and G347.3-0.5 and a single surviving continuous gravitational wave candidate

### v1

@@ Abstract
We carry out the deepest and broadest search for continuous gravitational-wave signals with frequencies between 20–1500 Hz, from three neutron stars at the center of the supernova remnants Cassiopeia A, Vela Jr., and G347.3–0.5. This search was made possible by the computing power shared by thousands of Einstein@Home volunteers. After the initial Einstein@Home search, we perform a multi-stage follow-up of the most promising $\approx$ 45 million signal candidates. In the last stages, we use independent data to further investigate the remaining candidates from the previous stages. We set the most stringent constraints to date on the gravitational-wave amplitude, equatorial ellipticity, r-mode saturation amplitude, and—for the first time—the neutron-star crustal anisotropy. For spin periods lower than 2 ms we constrain the ellipticity to be smaller than $4\times 10^{-7}$ for all targets. We exclude the crustal anisotropy to be smaller than $5\times 10^{-3}$ for spin periods between 1.3-100 ms. Only one candidate – from the low frequency G347.3 search – survives all follow-ups. We illustrate properties of this candidate. Investigations on new data will aid in clarifying its nature. Such “new” data exists and would be optimal for this purpose, but they are not publicly accessible at the time of writing.
@@ 1 Introduction
This manuscript is organized as follows. Section 2 introduces the targeted remnants; Section 3 describes the signal model for continuous gravitational waves; Section 4 describes the data sets that we use and their cleaning/conditioning; Section 5 describes the Einstein@Home search pipeline; Section 6 describes the hierarchical follow-up searches and their results; finally, Section 7 reports astrophysical constraints and discusses the astrophysical implications.
2 Targets
2.1 Cassiopeia A (G111.7-2.1)
@@ 4 The Data
5 The Einstein@Home Search
@@
The per‑segment coherent statistic $2\mathcal{F}_{i}$ combines data from both detectors within segment $i$, whereas the cross‑segment accumulation in Eq. 8 is incoherent. This mixed strategy is the reason the overall procedure is termed semi‑coherent. For computational efficiency, the initial evaluation of the detection statistics employs a fast approximation at each template. Candidates that rise to the top lists are then recomputed exactly at their nominal parameter points; we denote these refined values with the subscript “r”, e.g $2{\mathcal{F}_{r}}$ and $\hat{\beta}_{\mathrm{S/GLtL}r}$. True signals typically increase under this exact recomputation, while many noise outliers do not.
@@
We relax this model in two ways: i) we define the ranges of variability of $\dot{f}$ based on the range on $n$, and include in the possible $n$ range non-standard values, i.e. $n\rightarrow\infty$. ii) we further expand the searched waveforms to include ones for which the values of $\dot{f}$ and $\ddot{f}$ searched for a given $f$ are not consistent, in the sense that they correspond to different values of $n$. In particular for the upper bound on $\ddot{f}$ we take $|\dot{f}|=|\dot{f}|_{\textrm{max}}=f/\tau$, which from the first equation in 10 corresponds to $n=2$, and then further maximize setting $n=7$ in the second equation of Eqs. 10. This yields
@@
As the search ranges in $\dot{f}$ and $\ddot{f}$ expand with increasing $f$ (Eqs.11), the number of templates required in a given frequency band grows accordingly. The distribution of templates per Hz band for all targets is displayed in Fig.1. Although the parameter space of the high-frequency search is more than twice larger than that of the low-frequency search, a very similar number of templates are used in the low and high frequency bands. This is due to the fact that the search set-ups used in the low-frequency band are generally more sensitive than those in the high-frequency band. Similarly, the parameter space searched for Cas A is the largest, due to Cas A’s younger age, while the number of templates searched for Cas A is smaller compared to the other two targets, reflecting a less sensitive search. These are all consequences of the optimization scheme described in Ming2016, which, for a fixed computing budget, maximizes the overall detection probability by assigning the most sensitive set-ups to those frequency bands and targets that are most promising for a continuous wave detection.
@@
Table 2: The quantities $\Delta f$, $\Delta\dot{f}$, and $\Delta\ddot{f}$ denote, for each candidate, the one-sided search range used at Stages 1-4. For the initial Einstein@Home (E@H) search survey, the search ranges are full frequency range and its first two derivatives ranges as described in Eqs. 11. We use $N_{\mathrm{in}}$ for the number of candidates actually searched at Stages 1-4 (and for the total template count in the Einstein@Home run), and $N_{\mathrm{out}}$ for the number of survivors promoted from each stage. The quantity $R^{a}_{\mathrm{thr}}$ is the cut defined in Eq.13 (applied for $a=1,2,3,4$). Entries from the Einstein@Home survey are chosen via the pixeling scheme, so no $R_{\mathrm{thr}}$ is applied there. For G347.3 in the 20-500 Hz band, Stage-1 was not executed, because we skipped the Stage-1 follow-up search of it. For all three targets in the 500-1500 Hz band, Stage-4 was not executed, because we vetoed all candidates at Stage-3 and no more candidate needs to be verified at Stage-4.
6 Hierarchical follow up searches
@@
6.1 Stage-0 and candidate selection
@@
6.2 Coherence growth and target signal population
@@
The orientation angles are sampled uniformly, with $\cos\iota:[-1,1]$ and $\psi:[-\pi/4,\pi/4]$ (the latter reflecting the $\pi/2$ periodicity of the polarization angle’s fundamental domain, see Eqs.10 and 11 of JKS1998). Signal strengths are restricted to narrow intervals centered on the target $h_{0}$ values which are expected to yield approximately $90\%$ detection efficiency.
Less than 0.005% of the 45 million candidates have $2\bar{\mathcal{F}}_{r}^{\text{ Stage-0}}$ values lower than those of our test signal population. These candidates come from disturbed bands in the low-frequency region and they have extremely low $2\bar{\mathcal{F}}_{r}^{\text{ Stage-0}}$ values because of: (i) the trials factor due to frequency range: the number of templates searched per Hz in the low frequency regions is much smaller than at higher frequencies (see Figure 1 and Eqs. 11), so even weak candidates can reach the Einstein@Home top list. (ii) the trials factor due to top-list saturation: because the Einstein@Home results are pure top-lists (no pixeling), when a band is affected by a disturbance, most candidates come from the disturbed part of the 50-mHz band, leaving pixels outside the disturbed regions with very few candidates. Our candidates come from undisturbed pixels within a disturbed band, which means they are the top 5, picked among a very small set.
@@ 6.5 Stage-3 follow-ups
After Stage-3 all candidates from all the high frequency searches are rejected. The low-frequency Cas A search candidates are also all rejected. There only remains 1 candidate from the low frequency Vela Jr search and 13 candidates from the low frequency G347.3 search.
6.6 Stage-4 O4a follow-ups
@@ 6.7 The surviving candidate
Closer inspection of phase-parameter consistency between the results on O3 and O4a data is not convincing. We repeat Stage 3 and Stage 4 using the Bayesian follow-up method of Martins:2025jnq. We find that the O4 frequency posterior is not much different from its prior, which is the O3 posterior. If a third order frequency-derivative parameter is introduced all the O4 posteriors become more informative, but the overall O4 evidence decreases with respect to the O3 evidence.
@@
7 Results
7.1 Upper limits on the gravitational wave amplitude
@@
There are some half-Hz bands where no $h_{0}^{90\%}$ upper limit is quoted. As discussed in Section 4, frequency intervals dominated by prominent spectral lines are substituted with Gaussian noise. We apply the same substitution in the search-and–recovery Monte Carlo used to determine the upper limit, after the test signals have been added. This procedure may also remove parts of the signal. Depending on how much data are substituted, the resulting detection-efficiency curve may never attain the $90\%$ level, no matter how large the $h_{0}$ is. In such cases, we exclude these half-Hz bands from the upper limit statements, and add them to the excluded-band list, again available in machine readable format as the upper limit data. We exclude 7, 9, and 9 half-Hz bands between 20 and 500 Hz for Cas A, Vela Jr., and G347.3, respectively, and 42, 44, and 42 half-Hz bands between 500 and 1500 Hz for Cas A, Vela Jr., and G347.3, respectively.
@@ 7.1.1 Sensitivity Depth
Although the grid mismatch and $T_{\mathrm{coh}}$ we used for Cas A at low frequencies and Vela Jr. at high frequencies are the same, the sensitivity depth of the Cas A search is 6% higher than that of the Vela Jr search. The improvement is due to the fact that for Cas A at low frequencies, we followed up a larger fraction of candidates relative to the total search templates: The high-frequency Vela Jr. template bank has roughly 3 times more templates than the low-frequency Cas A search, yet only about 2 times as many candidates were followed up. This highlights that the achieved sensitivity depth depends not only on the initial search configuration (Table1) but also on post-processing choices—most notably, how many top-ranked candidates are followed-up.
@@
7.2 Constraints on Astrophysical Parameters
(a) 90 % upper limits on the ellipticity (a) and r-mode amplitude (b) of the two targets. For Vela Jr. we show two curves, corresponding to two distance estimates: 200 pc and 750 pc. For Cas A, we assume 3.4 kpc.
@@ 7.2 Constraints on Astrophysical Parameters
We can therefore recast the amplitude upper limits $h_{0}^{90\%}(f)$ as constraints on the equatorial ellipticity of the neutron star using Eq. 18. The corresponding curves are shown in Figure LABEL:fig:UL_e. It is expected that neutron star crusts can sustain ellipticities as large as $10^{-5}$, and young neutron stars may well present larger ellipticities than older neutron stars, acquired during the highly non-axysimmetric supernova event that generated them. Our searches probe a plausible range of ellipticities : smaller than $10^{-6}$ for all of the investigated remnants, for spin frequencies above 200 Hz. Particularly notable is the constraint for Vela Jr at 200 pc of $\epsilon\lesssim 4\times 10^{-8}$.
@@
Young neutron stars like our targets have long been identified as good candidates for r-mode gravitational wave emission, as they cool down after birth. Our upper limit for all targets exclude $\alpha\gtrsim 4\times 10^{-5}$ for spin frequencies smaller than 250 Hz, and for Vela Jr. at 200 pc, we can constrain $\alpha\lesssim 2\times 10^{-6}$. With predictions for the saturation amplitudes in the range $10^{-5}-10^{-3}$ (PhysRevD.76.064019; Haskell:2015iia), these upper limits are physically interesting.
@@
Since the anisotropy–induced ellipticity is frequency dependent, the corresponding braking index $n$ is also frequency dependent. Starting from the definition of the braking index, $n=\frac{\nu\ddot{\nu}}{\dot{\nu}^{2}}$, and combining this with Equation 20, we obtain
$n(f)=5+\frac{4\nu^{2}}{\nu^{2}-\nu_{0}^{2}}\,.$ (21)
The braking index $n(f)$ is shown on the right-hand $y$-axis of Figure 7. In the excluded frequency range, the effective braking index can reach values as high as 14.7 (blue vertical dashed line), which nevertheless remains within our searched parameter space because, as previously explained, the parameter space bounds of Eq. 11 are safely inclusive for all $n\geq 2$.

### v2

@@ Abstract
We carry out the deepest and broadest search for continuous gravitational-wave signals with frequencies between 20-1500 Hz, from three neutron stars at the center of the supernova remnants Cassiopeia A, Vela Jr., and G347.3-0.5. This search was made possible by the computing power shared by thousands of Einstein@Home volunteers. After the initial Einstein@Home search, we perform a multi-stage follow-up of the most promising $\approx$ 45 million signal candidates. In the last stages, we use independent data to further investigate the remaining candidates from the previous stages. We set the most stringent constraints to date on the gravitational-wave amplitude, equatorial ellipticity, r-mode saturation amplitude, and – for the first time – the neutron-star crustal anisotropy. For spin periods lower than 2 ms we constrain the ellipticity to be smaller than $4\times 10^{-7}$ for all targets. We exclude the crustal anisotropy to be smaller than $5\times 10^{-3}$ for spin periods between 1.3-100 ms. Only one candidate – from the low frequency G347.3 search – survives all follow-ups. We illustrate properties of this candidate. Investigations on new data will aid in clarifying its nature. Such “new” data exists and would be optimal for this purpose, but they are not publicly accessible at the time of writing. In the appendix we provide our estimate of the candidate parameters.
@@ I Introduction
This manuscript is organized as follows. Section II introduces the targeted remnants; Section III describes the signal model for continuous gravitational waves; Section IV describes the data sets that we use and their cleaning/conditioning; Section V describes the Einstein@Home search pipeline; Section VI describes the hierarchical follow-up searches and their results; finally, Section VII reports astrophysical constraints and discusses the astrophysical implications.
II Targets
II.1 Cassiopeia A (G111.7-2.1)
@@ IV The Data
V The Einstein@Home Search
@@
The per‑segment coherent statistic $2\mathcal{F}_{i}$ combines data from both detectors within segment $i$, whereas the cross‑segment accumulation in Eq. (8) is incoherent. This mixed strategy is the reason the overall procedure is termed semi‑coherent. For computational efficiency, the initial evaluation of the detection statistics employs a fast approximation at each template. Candidates that rise to the top lists are then recomputed exactly at their nominal parameter points; we denote these refined values with the subscript “r”, e.g $2{\mathcal{F}_{r}}$ and $\hat{\beta}_{\mathrm{S/GLtL}r}$. True signals typically increase under this exact recomputation, while many noise outliers do not.
@@
We relax this model in two ways: i) we define the ranges of variability of $\dot{f}$ based on the range on $n$, and include in the possible $n$ range non-standard values, i.e. $n\rightarrow\infty$. ii) we further expand the searched waveforms to include ones for which the values of $\dot{f}$ and $\ddot{f}$ searched for a given $f$ are not consistent, in the sense that they correspond to different values of $n$. In particular for the upper bound on $\ddot{f}$ we take $|\dot{f}|=|\dot{f}|_{\textrm{max}}=f/\tau$, which from the first equation in 10 corresponds to $n=2$, and then further maximize setting $n=7$ in the second equation of Eqs. (10). This yields
@@
As the search ranges in $\dot{f}$ and $\ddot{f}$ expand with increasing $f$ (Eqs. (11)), the number of templates required in a given frequency band grows accordingly. The distribution of templates per Hz band for all targets is displayed in Figure 1. Although the parameter space of the high-frequency search is more than twice larger than that of the low-frequency search, a very similar number of templates are used in the low and high frequency bands. This is due to the fact that the search set-ups used in the low-frequency band are generally more sensitive than those in the high-frequency band. Similarly, the parameter space searched for Cas A is the largest, due to Cas A’s younger age, while the number of templates searched for Cas A is smaller compared to the other two targets, reflecting a less sensitive search. These are all consequences of the optimization scheme described in Ming2016, which, for a fixed computing budget, maximizes the overall detection probability by assigning the most sensitive set-ups to those frequency bands and targets that are most promising for a continuous wave detection.
@@
Table 2: The quantities $\Delta f$, $\Delta\dot{f}$, and $\Delta\ddot{f}$ denote, for each candidate, the one-sided search range used at Stages 1-4. For the initial Einstein@Home (E@H) search survey, the search ranges are full frequency range and its first two derivatives ranges as described in Eqs. (11). We use $N_{\mathrm{in}}$ for the number of candidates actually searched at Stages 1-4 (and for the total template count in the Einstein@Home run), and $N_{\mathrm{out}}$ for the number of survivors promoted from each stage. The quantity $R^{a}_{\mathrm{thr}}$ is the cut defined in Eq. (13) (applied for $a=1,2,3,4$). Entries from the Einstein@Home survey are chosen via the pixeling scheme, so no $R_{\mathrm{thr}}$ is applied there. For G347.3 in the 20-500 Hz band, Stage-1 was not executed, because we skipped the Stage-1 follow-up search of it. For all three targets in the 500-1500 Hz band, Stage-4 was not executed, because we vetoed all candidates at Stage-3 and no more candidate needs to be verified at Stage-4.
VI Hierarchical follow up searches
@@
VI.1 Stage-0 and candidate selection
@@
VI.2 Coherence growth and target signal population
@@
The orientation angles are sampled uniformly, with $\cos\iota:[-1,1]$ and $\psi:[-\pi/4,\pi/4]$ (the latter reflecting the $\pi/2$ periodicity of the polarization angle’s fundamental domain, see Eqs. (10) and (11) of JKS1998). Signal strengths are restricted to narrow intervals centered on the target $h_{0}$ values which are expected to yield approximately $90\%$ detection efficiency.
Less than 0.005% of the 45 million candidates have $2\bar{\mathcal{F}}_{r}^{\text{ Stage-0}}$ values lower than those of our test signal population. These candidates come from disturbed bands in the low-frequency region and they have extremely low $2\bar{\mathcal{F}}_{r}^{\text{ Stage-0}}$ values because of: (i) the trials factor due to frequency range: the number of templates searched per Hz in the low frequency regions is much smaller than at higher frequencies (see Figure 1 and Eqs. (11)), so even weak candidates can reach the Einstein@Home top list. (ii) the trials factor due to top-list saturation: because the Einstein@Home results are pure top-lists (no pixeling), when a band is affected by a disturbance, most candidates come from the disturbed part of the 50-mHz band, leaving pixels outside the disturbed regions with very few candidates. Our candidates come from undisturbed pixels within a disturbed band, which means they are the top 5, picked among a very small set.
@@ VI.5 Stage-3 follow-ups
After Stage-3 all candidates from all the high frequency searches are rejected. The low-frequency Cas A search candidates are also all rejected. There only remains 1 candidate from the low frequency Vela Jr. search and 13 candidates from the low frequency G347.3 search.
VI.6 Stage-4 O4a follow-ups
@@ VI.7 The surviving candidate
Closer inspection of phase-parameter consistency between the results on O3 and O4a data is not convincing. We repeat Stage-3 and Stage-4 using the Bayesian follow-up method of Martins:2025jnq. We find that the O4 frequency posterior is not much different from its prior, which is the O3 posterior. If a third order frequency-derivative parameter is introduced all the O4 posteriors become more informative, but the overall O4 evidence decreases with respect to the O3 evidence.
@@
VII Results
VII.1 Upper limits on the gravitational wave amplitude
@@
There are some half-Hz bands where no $h_{0}^{90\%}$ upper limit is quoted. As discussed in Section IV, frequency intervals dominated by prominent spectral lines are substituted with Gaussian noise. We apply the same substitution in the search-and–recovery Monte Carlo used to determine the upper limit, after the test signals have been added. This procedure may also remove parts of the signal. Depending on how much data are substituted, the resulting detection-efficiency curve may never attain the $90\%$ level, no matter how large the $h_{0}$ is. In such cases, we exclude these half-Hz bands from the upper limit statements, and add them to the excluded-band list, again available in machine readable format as the upper limit data. We exclude 7, 9, and 9 half-Hz bands between 20 and 500 Hz for Cas A, Vela Jr., and G347.3, respectively, and 42, 44, and 42 half-Hz bands between 500 and 1500 Hz for Cas A, Vela Jr., and G347.3, respectively.
@@ VII.1.1 Sensitivity Depth
Although the grid mismatch and $T_{\mathrm{coh}}$ we used for Cas A at low frequencies and Vela Jr. at high frequencies are the same, the sensitivity depth of the Cas A search is 6% higher than that of the Vela Jr. search. The improvement is due to the fact that for Cas A at low frequencies, we followed up a larger fraction of candidates relative to the total search templates: The high-frequency Vela Jr. template bank has roughly 3 times more templates than the low-frequency Cas A search, yet only about 2 times as many candidates were followed up. This highlights that the achieved sensitivity depth depends not only on the initial search configuration (Table1) but also on post-processing choices—most notably, how many top-ranked candidates are followed-up.
@@
VII.2 Constraints on Astrophysical Parameters
(a) 90 % upper limits on the ellipticity (a) and r-mode amplitude (b) of the three targets as the function of gravitational wave signal frequency $f$ (twice of the spin frequency $\nu$). For Vela Jr. we show two curves, corresponding to two distance estimates: 200 pc and 750 pc. For Cas A, we assume 3.4 kpc.
@@ VII.2 Constraints on Astrophysical Parameters
We can therefore recast the amplitude upper limits $h_{0}^{90\%}(f)$ as constraints on the equatorial ellipticity of the neutron star using Eq. (18). The corresponding curves are shown in Figure LABEL:fig:UL_e. It is expected that neutron star crusts can sustain ellipticities as large as $10^{-5}$, and young neutron stars may well present larger ellipticities than older neutron stars, acquired during the highly non-axysimmetric supernova event that generated them. Our searches probe a plausible range of ellipticities : smaller than $10^{-6}$ for all of the investigated remnants, for spin frequencies above 200 Hz. Particularly notable is the constraint for Vela Jr. at 200 pc of $\varepsilon\lesssim 4\times 10^{-8}$.
@@
Young neutron stars like our targets have long been identified as good candidates for r-mode gravitational wave emission, as they cool down after birth. Our upper limit for all targets exclude $\alpha\gtrsim 4\times 10^{-5}$ for spin frequencies larger than 250 Hz, and for Vela Jr. at 200 pc, we can constrain $\alpha\lesssim 2\times 10^{-6}$. With predictions for the saturation amplitudes in the range $10^{-5}-10^{-3}$ (PhysRevD.76.064019; Haskell:2015iia), these upper limits are physically interesting.
@@
Since the anisotropy–induced ellipticity is frequency dependent, the corresponding braking index $n$ is also frequency dependent. Starting from the definition of the braking index, $n=\frac{\nu\ddot{\nu}}{\dot{\nu}^{2}}$, and combining this with Eq. (20), we obtain
$n(\nu)=5+\frac{4\nu^{2}}{\nu^{2}-\nu_{0}^{2}}\,.$ (21)
The braking index $n(\nu)$ is shown on the right-hand $y$-axis of Figure 7. In the excluded frequency range, the effective braking index can reach values as high as 14.7 (blue vertical dashed line), which nevertheless remains within our searched parameter space because, as previously explained, the parameter space bounds of Eq. (11) are safely inclusive for all $n\geq 2$.
@@ VIII Acknowledgments
Appendix A G347.3 Parameters
The best estimate for the G347.3 surviving candidate parameters is given in Table 3, from the posteriors of Figure 9. These parameters are based on the longest stretch of contiguous data available to us – O3a+b – and leverage the Bayesian follow-up method of Martins:2025jnq.
$f$ [Hz] $\dot{f}$ [Hz/s] $\ddot{f}$ [Hz/s2] $\Delta^{99\%}f$ [Hz] $\Delta^{99\%}\dot{f}$ [Hz/s] $\Delta^{99\%}\ddot{f}$ [Hz/s2]
$31.691150039$ $3.598467\times 10^{-10}$ $7.744\times 10^{-20}$ $\pm 1.5\times 10^{-8}$ $\pm 5.0\times 10^{-15}$ $\pm 8.0\times 10^{-22}$
Table 3: Maximum likelihood estimator values and 99% credible intervals from the O3a+b follow-up of the surviving G347.3 candidate.
Figure 9: O3a+b Bayesian follow-up posteriors and credible intervals around the maximum likelihood estimator.